OpenStudy (darth_vader):

The sum of three consecutive even integers is 72. Let x represent the least consecutive integer. What equation could be used to find the least consecutive integer? What is the second consecutive integer? Enter your answers in the boxes. Equation: Second consecutive integer:

OpenStudy (benji):

Alright I will give you the equation, then you tell me what the three integers are okay? \[72=x+x+1+x+2\] or \[72=3x+3\]

OpenStudy (darth_vader):

2 and 1?

OpenStudy (benji):

Find x.

OpenStudy (darth_vader):

sorry but i aint good at math that well but i will try at lest

OpenStudy (benji):

Okay\[72=3x+3\] \[72-3=3x+3-3\] \[69=3x\] \[\frac{ 69 }{ 3 }=\frac{ 3x }{ 3 }\] \[x=?\]

OpenStudy (darth_vader):

32?

OpenStudy (benji):

69/3 = 23

OpenStudy (benji):

So we found x which is the least consecutive number. So the other 2 are?

OpenStudy (darth_vader):

??????????

OpenStudy (benji):

x = 23 so what are x+1 and x+2?

OpenStudy (darth_vader):

hmmmm

OpenStudy (darth_vader):

idk

OpenStudy (benji):

23+1 and 23+2, all three consecutive (which means right after eachother) is 23, 24, and 25. 23+24+25=72

OpenStudy (mathmale):

Darth: Please show your work, not just spout out possible answers. If the first of these 3 consecutive integers is x, how would you represent the second? The third? It is critically important that you understand how to represent unknowns algebraically.